OpenStudy (anonymous):

Hi coud you guys help me solving this trig equation::: cos^2x + 3 sin2x - 3 = 0 , find all the values that satisfies the equation.

OpenStudy (anonymous):

this is what is have so far i ahve tried using some identities,

OpenStudy (freckles):

By the way, I'm here thinking about your problem.

OpenStudy (anonymous):

thank you.

OpenStudy (freckles):

So this one looks pretty tricky.

OpenStudy (anonymous):

yeah it's really hard

OpenStudy (freckles):

I think I might have something...

OpenStudy (freckles):

there was a mistake rethinking

OpenStudy (freckles):

This is just a question from the trig book?

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

@aum any ideas ?

OpenStudy (anonymous):

how about diving by cos^2x

OpenStudy (freckles):

That maybe getting rid of some our solutions thought because cos can be 0 sometimes and we can't divide an equation by 0

OpenStudy (anonymous):

yeah i was told that

OpenStudy (anonymous):

hmmmm

OpenStudy (freckles):

@ganeshie8 what do you think here?

OpenStudy (freckles):

what book does this come from

OpenStudy (freckles):

also make sure there is no type-o

OpenStudy (anonymous):

hmm the exersice is from a homework, the teacher gave it to us actually but dont really know what textbook he got that from

OpenStudy (freckles):

I honestly feel like this question cannot be solved by just trigonometric identities If the equation was \[ cos^2(2x)+2sin(2x)-3=0\] but we could easily solve it

OpenStudy (freckles):

\[1-\sin^2(2x)+2\sin(2x)-3=0 \] \[-\sin^2(2x)+2\sin(2x)-2=0\]

OpenStudy (freckles):

of if that one two was three whatever like we would then use the quadratic formula...

OpenStudy (freckles):

even wolfram gives answers not in the exact form for the equation you gave and this could mean either we can't solve it just be trig equations or that there is a very creative way to find the exact solutions

OpenStudy (anonymous):

ohh

OpenStudy (freckles):

Yes I had to cheat and look at wolfram All of those alternate forms are very ugly They don't one pretty alternate form listed

OpenStudy (anonymous):

i got this using graph

OpenStudy (freckles):

You could give approximate solutions :( But I don't see how to get exact solutions

OpenStudy (anonymous):

hmm i made a mistake this one looks like on wolfram

OpenStudy (anonymous):

ohh no way ho w do i solve that thing

OpenStudy (freckles):

i will try something weird make a right triangle  let u=sin(2x) so cos^2(x)=(cos(2x)+1)/2 so we have cos^2(x)+3sin(2x)-3=0 \[\frac{\sqrt{1-u^2}+1}{2}+3u-3=0\]